OpenStudy (anonymous):

In how many different ways can the top nine new indie bands be ranked on a top nine list? The top hit song for each of the nine bands will compete to receive monetary awards of $1000, $500, $250 and $100, respectively. In how many ways can the awards be given out? Show steps please.

OpenStudy (anonymous):

must be on line class. seen this one before. "indie bands" how au courant.

OpenStudy (anonymous):

first one is 9!

OpenStudy (anonymous):

second one is \[9\times 8\times 7\times 6\]

OpenStudy (anonymous):

9!=362880

OpenStudy (anonymous):

Yes, it's homework help for a online class ..Thanks so much..I ran out of tutor hours and have to wait until tom. to get some homework help from a tutor and this is due today....